OpenStudy (anonymous):

Solve for x, given that EF is the median of trapezoid ABCD. Find AD and EF BC=12 EF=5x-20 AD=4x-10?

OpenStudy (anonymous):

median =(base1+base2)/2 5x-20=(12+4x-10)/2 10x-40=12+4x-10 6x=42 x=7 EF=5(7)-20=15 AD=4(7)-10=18

OpenStudy (anonymous):

thank u so much! @AJ01